A flat surface of area 2.60 m2 is rotated in a uniform electric field of magnitude E-5.90 x 105 N/C. (a) Determine the electric flux through this area when the electric field is perpendicular to the surface. The correct answer is not zero. N m2/C (b) Determine the electric flux through this area when the electric field is parallel to the surface. 15 34 Your response differs significantly from the correct answer. Rework your solution from the beginning and check each step carefully. N·m2/C
Answer #1

Electric Flux, F = E*dA*cos theta

a) When electric field is perpendicular to the surface,

Angle between electric field E and dA is zero

F = 5.90*10^5 N/C*2.60 m^2*cos 0 degree

= 1.534*10^6 N m^2/C

b) When electric field is parallel to the surface

Angle between electric field E and dA is 90 degree

F = 5.90*10^5*2.60*cos 90 degree

= 0